# Songrun He   Linying Lv   Asaf Manela   Jimmy Wu

### v2

@@
(First draft: February 2025. This draft: March 2025.)
Large language models are increasingly used in social sciences, but their training data can introduce lookahead bias and training leakage. A good chronologically consistent language model requires efficient use of training data to maintain accuracy despite time-restricted data. Here, we overcome this challenge by training a suite of chronologically consistent large language models, ChronoBERT and ChronoGPT, which incorporate only the text data that would have been available at each point in time. Despite this strict temporal constraint, our models achieve strong performance on natural language processing benchmarks, outperforming or matching widely used models (e.g., BERT), and remain competitive with larger open-weight models. Lookahead bias is model and application-specific because even if a chronologically consistent language model has poorer language comprehension, a regression or prediction model applied on top of the language model can compensate. In an asset pricing application predicting next-day stock returns from financial news, we find that ChronoBERT’s real-time outputs achieve a Sharpe ratio comparable to state-of-the-art models, indicating that lookahead bias is modest. Our results demonstrate a scalable, practical framework to mitigate training leakage, ensuring more credible backtests and predictions across finance and other social science domains.
@@ 1 Introduction
We address this challenge by training chronologically consistent LLMs trained exclusively on historical textual data available at the time. While training language models with historical data timestamped at the point of its availability is conceptually straightforward, ensuring these models are competitive with state-of-the-art counterparts remains a significant challenge.
@@
Our chief contribution is ChronoBERT and ChronoGPT: a pair of models for each year starting in 1999 that never see future data during training. On the GLUE benchmark (Wang et al., 2019), even our earliest ChronoBERT outperforms existing no-leakage models, including StoriesLM (Sarkar and Vafa, 2024) and FinBERT (Huang et al., 2023). It also matches or surpasses the popular BERT (Devlin et al., 2019), which ranks first in downloads among all language models on Hugging Face as of February 2025.111Based on download statistics from Hugging Face at https://huggingface.co/models?sort=downloads. Strong language understanding, we show, does not require data from the future.
On the HellaSwag sentence-completion task (Zellers et al., 2019), all ChronoGPT vintages surpass the language understanding ability of OpenAI’s GPT of similar size. Thus ChronoBERT and ChronoGPT emerge as the strongest leak-free LLMs to date. They also maintain a modest scale, making them feasible for large embedding tasks where massive models (e.g., Llama) are too costly.
@@
We test for lookahead bias by forecasting stock returns from news with ChronoBERT and ChronoGPT. Leveraging extensive financial newswire data, we find that the portfolio performance of ChronoBERT and ChronoGPT matches that of the state-of-the-art Llama 3.1 (Dubey et al., 2024), with both models delivering economically substantial and statistically significant gains compared to StoriesLM (Sarkar, 2024) and FinBERT (Huang et al., 2023). Our findings suggest that lookahead bias in this context is relatively modest.
@@
Our contribution is threefold. First, we explain how to build chronologically consistent LLMs that preserve validity in economic and financial tests. Second, we find that strong language understanding can be attained without introducing training leakage: ChronoBERT and ChronoGPT remain competitive with open-weight baselines. Third, we measure lookahead bias in news-based return forecasting and find it modest. These points open a path to more credible LLM applications across the social sciences.
@@
A second strand of literature pertains to the application and development of natural language processing (NLP) tools for financial economics research (Hoberg and Manela, 2025). Methodological advancements in this area have evolved from early dictionary-based approaches (Tetlock, 2007; Loughran and McDonald, 2011), to text regressions (Manela and Moreira, 2017; Kelly et al., 2021), to topic modeling (Bybee et al., 2024), and most recently, to the integration of LLMs (Jha et al., 2025; Chen et al., 2023; Lv, 2024). These developments underscore the growing demand for more advanced and scalable tools to answer research questions in finance and economics.
In this broader context, our main contribution is a framework for developing LLMs that are both free of lookahead bias and capable of high-level language comprehension. Our approach does not require masking (Glasserman and Lin, 2023), which can destroy information. Instead, we train a series of chronologically consistent models with knowledge cutoffs from 1999 to 2024 which offer superior language understanding and more up-to-date knowledge to similar such attempts (e.g., Sarkar, 2024). [^2] In contemporaneous work, Rahimikia and Drinkall (2024) introduce a series of LLMs trained on time-stamped proprietary financial texts for predictions. They do not evaluate the models’ general language understanding capabilities, but based on the financial forecasting metrics they do provide, their best-performing model generates a long-short portfolio Sharpe ratio of 3.45 from 2017 to 2023, significantly underperforming ChronoBERT, which achieves an SR of 4.50 over the same period. Unfortunately, the authors have since retracted their models from Hugging Face, preventing further direct comparison.
@@ 2.1 Pretraining Methodology for ChronoBERT
When pretraining ChronoBERT, we incorporate a state-of-the-art BERT architecture from Portes et al. (2023) and Warner et al. (2024). [^3] We thank the authors for providing their pretraining code at https://github.com/mosaicml/examples/tree/main/examples/benchmarks/bert and https://github.com/AnswerDotAI/ModernBERT. Compared to the original BERT model by Devlin et al. (2019), this enhanced architecture integrates recent advancements in rotary positional embeddings that support longer context lengths and employ flash attention, significantly improving pretraining efficiency and computational speed.
For the pretraining task, we follow Warner et al. (2024) by adopting masked token prediction while omitting the next sequence prediction task, as prior research has shown the latter increases training overhead without meaningful performance gains. [^4] We provide details on ChronoBERT’s masked language modeling pretraining objective and attention mechanisms in Appendix B and contrast them with those of ChronoGPT.
@@ 2.2 Pretraining Methodology for ChronoGPT
ChronoGPT is trained using a causal language modeling (CLM) objective, learning to predict the next token based on previously observed sequences. [^7] We provide details on ChronoGPT’s causal language modeling pretraining objective and attention mechanisms in Appendix B and contrast them with those of ChronoBERT. Given the efficiency gains provided by the modded-nanoGPT optimizer, ChronoGPT requires fewer training iterations to reach competitive performance. We adopt a low-epoch, high-throughput strategy, allowing the model to rapidly extract essential linguistic and contextual structures. Our first model checkpoint, $\text{ChronoGPT}_{1999}$, is trained on 21 billion tokens over approximately 3 epochs, striking a balance between computational efficiency and representational robustness.
@@ 2.3.1 Language Understanding
The GLUE evaluation framework, introduced by Wang et al. (2019), comprises multiple classification tasks designed to measure a model’s language understanding. [^8] The details and leaderboard of GLUE evaluation can be found at https://gluebenchmark.com/. Appendix A provides an overview of the eight GLUE tasks. Appendix B describes the architectural differences between ChronoBERT and ChronoGPT for sequence classification in GLUE evaluation. This framework was also the primary evaluation metric used in Devlin et al. (2019) to assess BERT’s language capabilities.
@@
Beyond GLUE, we evaluate ChronoGPT’s commonsense reasoning with HellaSwag (Zellers et al., 2019) in a zero-shot setting. HellaSwag presents a sentence completion task where each context ends with an incomplete phrase (e.g., “A woman sits at a piano,”) and the model must choose the most plausible continuation from four possible endings (e.g., “She sets her fingers on the keys.”). We compute the likelihood of each candidate ending under ChronoGPT’s autoregressive predictions and select the most likely continuation. Because HellaSwag employs adversarial filtering to create highly deceptive distractors, strong performance on this task signals robust contextual and procedural reasoning. [^9] The leaderboard and data of HellaSwag can be found at https://paperswithcode.com/sota/sentence-completion-on-hellaswag. Appendix A provides detailed information on the methodology used for HellaSwag evaluations.
@@ 3.1 Pretraining Model Fit
These improvements translate into enhanced language understanding, as reflected in the GLUE and HellaSwag scores. ChronoBERT begins outperforming BERT after approximately 350 billion training tokens and continues to improve thereafter, while ChronoGPT surpasses a GPT-2 model of the same size after 21 billion training tokens. [^14] We obtained the Hellaswag evaluation score for the GPT‑2 model from Karpathy (2024), available at https://github.com/karpathy/build-nanogpt/blob/master/play.ipynb.
@@
We find that with the introduction of new data, the validation loss continues to drop. Starting in the year 2013, we introduce high-quality common crawl data. We witness a significant decrease in validation loss with the increase in data diversity. In the right panel of Figure 2, the GLUE and HellaSwag scores for nearly all models exceed that of BERT and GPT-2, highlighting their superior language understanding and overall quality.
@@
The left panel shows the validation loss, measured using cross-entropy loss and language prediction accuracy, as the $\text{ChronoBERT}_{1999}$ and $\text{ChronoGPT}_{1999}$ are trained on an increasing number of tokens. The right panel displays the GLUE scores and HellaSwag values as training progresses. For ChronoBERT and ChronoGPT, the final model checkpoint is trained on 460 billion tokens and 21 billion tokens, respectively. The training corpus consists of text up to December 1999.
@@ 3.2 Language Understanding
StoriesLM: A pretrained model from Sarkar and Vafa (2024), trained on historical news articles. We use the final version trained on data up to 1963.171717Model downloaded from https://huggingface.co/StoriesLM/StoriesLM-v1-1963.
GPT-2: The 124 million parameter version of OpenAI’s GPT-2 model from Radford et al. (2019). [^18] Model downloaded from https://huggingface.co/openai-community/gpt2.
@@
$\text{ChronoGPT}_{1999}$ 124M 1,792 December, 1999
$\text{ChronoGPT}_{2024}$ 124M 1,792 December, 2024
@@
COLA 57.32 56.32 37.13 31.70
SST2 91.82 92.58 89.68 88.53
MRPC 92.71 92.45 82.92 85.34
STSB 89.57 89.93 81.57 82.58
QQP 88.54 88.90 82.43 83.53
MNLI 86.19 86.89 77.63 79.15
QNLI 90.61 92.04 84.94 85.98
RTE 80.94 85.20 67.08 67.80
GLUE 84.71 85.54 75.42 75.58
@@
By contrast, $\text{ChronoGPT}_{2024}$ and $\text{ChronoGPT}_{1999}$ do not match ChronoBERT’s scores on GLUE. Their lower performance reflects the nature of an autoregressive (decoder-only) architecture, which processes text strictly left to right for text generation. While this approach can excel at generative tasks, it often underperforms on classification benchmarks that benefit from bidirectional processing of the entire input.
Overall, ChronoBERT exhibits strong language understanding, coming remarkably close to Llama 3.1’s performance despite Llama 3.1 being a much larger and more recent model. The performance gap between ChronoBERT and StoriesLM likely stems from differences in training scale (460B versus 19B tokens) and the quality and diversity of the training data (ChronoBERT’s high-quality corpus versus StoriesLM’s unfiltered news-only dataset). Similarly, ChronoBERT’s significant edge over FinBERT highlights the importance of diverse and high-quality pretraining data, as FinBERT’s domain-specific financial texts lack comprehensive quality checks.
Importantly, our chronologically consistent models also match or exceed the performance of BERT and GPT-2 despite their chronologically bounded pretraining data, demonstrating that high-quality LLMs can be constructed without chronological leakage. This makes them particularly valuable for applications that require good language understanding and no lookahead bias. By maintaining both strong language understanding and temporal integrity, our models are particularly well-suited for applications where preserving the timeline of information is critical, making it a robust choice for a wide range of downstream tasks.
@@ 3.3 Validation of Chronological Consistency
Prompt year: 1992 2000 2008 2016 2020 2024
BERT Clinton Clinton Obama Obama Obama Obama
$\text{ChronoBERT}_{1999}$ Roosevelt Roosevelt Roosevelt Hoover Hoover Washington
$\text{ChronoBERT}_{2000}$ Clinton Clinton Clinton Clinton Clinton Wilson
$\text{ChronoBERT}_{2001}$ Clinton Clinton Clinton Clinton Clinton Washington
$\text{ChronoBERT}_{2002}$ Clinton Bush Bush Clinton Bush Clinton
$\text{ChronoBERT}_{2003}$ Clinton Bush Bush Clinton Bush Clinton
$\text{ChronoBERT}_{2004}$ Clinton Bush Bush Clinton Bush Clinton
$\text{ChronoBERT}_{2005}$ Clinton Bush Bush Bush Bush Clinton
$\text{ChronoBERT}_{2006}$ Clinton Bush Bush Bush Bush Clinton
$\text{ChronoBERT}_{2007}$ Clinton Bush Bush Bush Bush Monroe
$\text{ChronoBERT}_{2008}$ Clinton Bush Bush Obama Bush Wilson
$\text{ChronoBERT}_{2009}$ Clinton Clinton Obama Obama Obama Wilson
$\text{ChronoBERT}_{2010}$ Clinton Obama Obama Obama Obama Wilson
$\text{ChronoBERT}_{2011}$ Clinton Clinton Obama Obama Obama Wilson
$\text{ChronoBERT}_{2012}$ Obama Obama Obama Obama Obama Obama
$\text{ChronoBERT}_{2013}$ Clinton Obama Obama Obama Obama Monroe
$\text{ChronoBERT}_{2014}$ Clinton Bush Obama Obama Obama Monroe
$\text{ChronoBERT}_{2015}$ Clinton Clinton Obama Obama Obama Monroe
$\text{ChronoBERT}_{2016}$ Clinton Bush Obama Obama Obama Obama
$\text{ChronoBERT}_{2017}$ Clinton Bush Obama Trump Trump Monroe
$\text{ChronoBERT}_{2018}$ Clinton Bush Obama Trump Trump Obama
$\text{ChronoBERT}_{2019}$ Clinton Bush Obama Trump Trump Obama
$\text{ChronoBERT}_{2020}$ Clinton Bush Obama Trump Trump Trump
$\text{ChronoBERT}_{2021}$ Clinton Clinton Obama Trump Biden Biden
$\text{ChronoBERT}_{2022}$ Clinton Bush Obama Trump Biden Biden
$\text{ChronoBERT}_{2023}$ Clinton Bush Obama Trump Biden Biden
$\text{ChronoBERT}_{2024}$ Clinton Bush Obama Trump Biden Biden
Table 3: Mask Language Modeling-Based Predictions of U.S. Presidents
This table displays ChronoBERT’s masked-token predictions to the sentence “After the {year} U.S. presidential election, President [MASK] was inaugurated as U.S. President in the year {year+1}.” The gray area denotes predictions covering the post-knowledge cutoff period, including election years where the president-elect has yet to be inaugurated. The blue text highlights correct predictions. For reference, predictions from BERT (released in October 2018) are also included.
@@
To detect leakage in the textual data used to pretrain our chronologically consistent models, we evaluate them on events occurring after the model’s knowledge cutoff. For ChronoBERT, this involves constructing a sentence with a masked token representing time-specific information beyond the model’s cutoff. We use each model to tokenize the sentence and predict the masked token. For example, consider:

### v3

@@
(First draft: February 2025. This draft: July 2025.)
Large language models are increasingly used in social sciences, but their training data can introduce lookahead bias and training leakage. A good chronologically consistent language model requires efficient use of training data to maintain accuracy despite time-restricted data. Here, we overcome this challenge by training a suite of chronologically consistent large language models, ChronoBERT and ChronoGPT, which incorporate only the text data that would have been available at each point in time. Despite this strict temporal constraint, our models achieve strong performance on natural language processing benchmarks, outperforming or matching widely used models (e.g., BERT), and remain competitive with larger open-weight models. Lookahead bias is model and application-specific because even if a chronologically consistent language model has poorer language comprehension, a regression or prediction model applied on top of the language model can compensate. In an asset pricing application predicting next-day stock returns from financial news, we find that ChronoBERT and ChronoGPT’s real-time outputs achieve Sharpe ratios comparable to a much larger Llama model, indicating that lookahead bias is modest. Our results demonstrate a scalable, practical framework to mitigate training leakage, ensuring more credible backtests and predictions across finance and other social science domains.
@@ 1 Introduction
We address this challenge by training chronologically consistent LLMs trained exclusively on historical textual data available at the time. While training language models with historical data timestamped at the point of its availability is conceptually straightforward, ensuring these models are competitive with highly capable open-weight counterparts remains a significant challenge.
@@
Our chief contribution is ChronoBERT and ChronoGPT: a pair of models for each year starting in 1999 that never see future data during training. On the GLUE benchmark (Wang et al., 2019), Both ChronoBERT and ChronoGPT outperform existing no-leakage models, including StoriesLM (Sarkar, 2024) and FinBERT (Huang et al., 2023). Even our earliest ChronoBERT matches or surpasses the popular BERT (Devlin et al., 2019), which ranks first in downloads among all language models on Hugging Face as of February 2025.111Based on download statistics from Hugging Face at https://huggingface.co/models?sort=downloads. Strong language understanding, we show, does not require data from the future. While ChronoBERT and ChronoGPT emerge as the strongest leak-free LLMs to date, they also maintain a modest scale, making them feasible for large embedding tasks where massive models (e.g., Llama) are too costly.
@@
We test for lookahead bias by forecasting stock returns from news with ChronoBERT and ChronoGPT. Leveraging extensive financial newswire data, we find that the portfolio performance of ChronoBERT and ChronoGPT matches that of Llama 3.1 (Dubey et al., 2024), with both models delivering economically substantial and statistically significant gains compared to StoriesLM (Sarkar, 2024) and FinBERT (Huang et al., 2023). Our findings suggest that lookahead bias in this context is relatively modest.
@@
Our contribution is threefold. First, we explain how to build chronologically consistent LLMs that preserve validity in economic and financial tests. Second, we find that strong language understanding can be attained without introducing training leakage: ChronoBERT and ChronoGPT remain competitive with powerful, open-weight baselines. Third, we measure lookahead bias in news-based return forecasting and find it modest. These points open a path to more credible LLM applications across the social sciences.
@@
A second strand of literature pertains to the application and development of natural language processing (NLP) tools for financial economics research (Hoberg and Manela, 2025). Methodological advancements in this area have evolved from early dictionary-based approaches (Tetlock, 2007; Loughran and McDonald, 2011), to text regressions (Manela and Moreira, 2017; Kelly et al., 2021), to topic modeling (Bybee et al., 2024), to neural embedding methods (Goetzmann et al., 2022), and most recently, to the integration of LLMs (Jha et al., 2025; Chen et al., 2023; Lv, 2024). These developments underscore the growing demand for more advanced and scalable tools to answer research questions in finance and economics.
In this broader context, our main contribution is a framework for developing LLMs that are both free of lookahead bias and capable of high-level language comprehension. Our approach does not require masking (Glasserman and Lin, 2023; Engelberg et al., 2025), which can destroy information, and may not completely remove the look-ahead bias of high-memory LLMs (Levy, 2024; Lopez-Lira et al., 2025; Didisheim et al., 2025). Instead, we train a series of chronologically consistent models with knowledge cutoffs from 1999 to 2024 that offer superior language understanding and more up-to-date knowledge than similar attempts (e.g., Sarkar, 2024). [^2] In contemporaneous work, Rahimikia and Drinkall (2024) introduce a series of LLMs trained on time-stamped proprietary financial texts for predictions. They do not evaluate the models’ general language understanding capabilities, but based on the financial forecasting metrics they do provide, their best-performing model generates a long-short portfolio Sharpe ratio of 3.45 from 2017 to 2023, significantly underperforming ChronoGPT and ChronoBERT. Unfortunately, the authors have since retracted their models from Hugging Face, preventing further direct comparison.
@@ 2.1 Pretraining Methodology for ChronoBERT
When pretraining ChronoBERT, we incorporate a modern BERT architecture from Portes et al. (2023) and Warner et al. (2024). [^3] We thank the authors for providing their pretraining code at https://github.com/mosaicml/examples/tree/main/examples/benchmarks/bert and https://github.com/AnswerDotAI/ModernBERT. Compared to the original BERT model by Devlin et al. (2019), this enhanced architecture integrates recent advancements in rotary positional embeddings that support longer context lengths and employ flash attention, significantly improving pretraining efficiency and computational speed.
For the pretraining task, we follow Warner et al. (2024) by adopting masked token prediction while omitting the next sequence prediction task, as prior research has shown the latter increases training overhead without meaningful performance gains. [^4] We provide details on ChronoBERT’s masked language modeling pretraining objective and attention mechanisms in Appendix C and contrast them with those of ChronoGPT.
@@ 2.2 Pretraining Methodology for ChronoGPT
ChronoGPT is trained using a causal language modeling (CLM) objective, learning to predict the next token based on previously observed sequences. [^7] We provide details on ChronoGPT’s causal language modeling pretraining objective and attention mechanisms in Appendix C and contrast them with those of ChronoBERT.
To determine the optimal model size, we first conducted a scaling law experiment to assess the impact of parameter count within our data-constrained, chronological training regime. As detailed in Appendix B, this test revealed that increasing model size from 124M to 1.5B consistently improved performance on both validation loss and language understanding. Based on these findings, we selected the 1.5B parameter model, as it offered the best performance while remaining within our computational budget.
Given the efficiency gains provided by the modded-nanoGPT optimizer, ChronoGPT requires fewer training iterations to reach competitive performance. We adopt a low-epoch, high-throughput strategy, allowing the model to rapidly extract essential linguistic and contextual structures. Our first vintage model, $\text{ChronoGPT}_{1999}$, is trained on 71 billion tokens over approximately 10 epochs, striking a balance between computational efficiency and representational robustness.
@@ 2.3.1 Language Understanding
The GLUE evaluation framework, introduced by Wang et al. (2019), comprises multiple classification tasks designed to measure a model’s language understanding. [^8] The details and leaderboard of GLUE evaluation can be found at https://gluebenchmark.com/. Appendix A provides an overview of the eight GLUE tasks. Appendix C describes the architectural differences between ChronoBERT and ChronoGPT for sequence classification in GLUE evaluation. This framework was also the primary evaluation metric used in Devlin et al. (2019) to assess BERT’s language capabilities.
@@
Beyond GLUE, we evaluate ChronoGPT’s commonsense reasoning with HellaSwag (Zellers et al., 2019) in a zero-shot setting. HellaSwag presents a sentence completion task where each context ends with an incomplete phrase (e.g., “A woman sits at a piano,”) and the model must choose the most plausible continuation from four possible endings (e.g., “She sets her fingers on the keys.”). We compute the likelihood of each candidate ending under ChronoGPT’s autoregressive predictions and select the most likely continuation. Since HellaSwag employs adversarial filtering to create highly deceptive distractors, strong performance on this task signals robust contextual and procedural reasoning. [^9] The leaderboard and data of HellaSwag can be found at https://paperswithcode.com/sota/sentence-completion-on-hellaswag. Appendix A provides detailed information on the methodology used for HellaSwag evaluations.
@@ 3.1 Pretraining Model Fit
These improvements translate into enhanced language understanding, as reflected in the GLUE and HellaSwag scores. ChronoBERT surpasses the original BERT’s performance on the GLUE benchmark after being trained on approximately 350 billion tokens. Similarly, ChronoGPT outperforms GPT-2 on the HellaSwag benchmark after around 3 billion training tokens.
When evaluated against more demanding benchmarks, the performance of our models begins to plateau. We compare ChronoBERT to ModernBERT, which is trained on a significantly larger text corpus, and ChronoGPT to GPT-2 XL, which has a comparable number of parameters. ChronoBERT’s GLUE score stabilizes at 85, falling short of ModernBERT’s 88. Likewise, ChronoGPT achieves a HellaSwag score of 37% after 40 billion tokens, which is lower than the 48% achieved by GPT-2 XL. We attribute this performance ceiling to the temporally-constrained nature of our training data, as continued training on the same dataset yields no further improvements.
@@
We find that with the introduction of new data, the validation loss continues to drop. Starting in the year 2013, we introduce high-quality common crawl data. We witness a significant decrease in validation loss with the increase in data diversity. In the right panel of Figure 2, the GLUE and HellaSwag scores for nearly all models exceed those of BERT and GPT-2, highlighting their superior language understanding and overall quality. [^14] ModernBERT is trained on roughly 1.8 trillion tokens, over three times the 525 billion tokens used for $\text{ChronoBERT}_{2024}$. This much larger pre-training corpus largely explains its stronger GLUE performance.
@@
The left panel shows the validation loss, measured using cross-entropy loss and language prediction accuracy, as the $\text{ChronoBERT}_{1999}$ and $\text{ChronoGPT}_{1999}$ are trained on an increasing number of tokens. The right panel displays the GLUE scores and HellaSwag values as training progresses. For ChronoBERT and ChronoGPT, the final model checkpoint is trained on 460 billion tokens and 70 billion tokens, respectively. The training corpus consists of text up to December 1999.
@@ 3.2 Language Understanding
StoriesLM: A pretrained model from Sarkar (2024), trained on historical news articles. We use the final version trained on data up to 1963.181818Model downloaded from https://huggingface.co/StoriesLM/StoriesLM-v1-1963.
@@
$\text{ChronoGPT}_{1999}$ 1,552M 1,792 December, 1999
$\text{ChronoGPT}_{2024}$ 1,552M 1,792 December, 2024
Llama 3.1 8,030M 128,000 December 2023
@@
COLA 57.32 56.32 35.50 48.20
SST2 91.82 92.58 90.60 92.40
MRPC 92.71 92.45 87.66 88.55
STSB 89.57 89.93 80.65 86.07
QQP 88.54 88.90 83.70 85.75
MNLI 86.19 86.89 82.27 84.91
QNLI 90.61 92.04 84.68 88.56
RTE 80.94 85.20 74.30 78.63
GLUE 84.71 85.54 77.42 81.63
@@
Similarly, the ChronoGPT models showcase the effectiveness of our training methodology, with $\text{ChronoGPT}_{2024}$ (81.63) significantly outperforming FinBERT and StoriesLM. The most salient finding for ChronoGPT is its growth; it gains over 4 points on the GLUE benchmark from its 1999 counterpart (77.42), with especially large improvements in grammatical correctness (COLA) and inference (RTE). This confirms that chronological data enrichment progressively enhances language understanding in both encoder-only and decoder-only architectures.
Overall, ChronoBERT and ChronoGPT exhibit strong language understanding, with ChronoBERT coming remarkably close to Llama 3.1’s performance despite Llama 3.1 being a much larger and more recent model. The performance gap between ChronoBERT and StoriesLM likely stems from differences in training scale (460B versus 19B tokens) and the quality and diversity of the training data (ChronoBERT’s high-quality corpus versus StoriesLM’s unfiltered news-only dataset). Similarly, ChronoBERT and ChronoGPT’s significant edge over FinBERT highlights the importance of diverse and high-quality pretraining data, as FinBERT’s domain-specific financial texts lack comprehensive quality checks.
Importantly, our chronologically consistent models prove that lookahead bias can be eliminated without compromising the core capabilities expected of modern LLMs. This makes them particularly valuable for applications that require good language understanding and no lookahead bias. By maintaining both strong language understanding and temporal integrity, our models are particularly well-suited for applications where preserving the timeline of information is critical, making them a robust choice for a wide range of downstream tasks.
@@ 3.3 Validation of Chronological Consistency
Although we exercise great caution in curating our training data to ensure they include only information that we believe was available by a specific date, this process is not necessarily without imperfections. Potential inaccuracies in recorded publication dates—for example, if a hard copy document was digitized via optical character recognition (OCR) and an incorrect date was captured in the process—may result in the unintended inclusion of information that was not actually available at the intended time, thereby introducing traces of lookahead bias into our models.
@@
To detect leakage in the textual data used to pretrain our chronologically consistent models, we evaluate them on events occurring after the model’s knowledge cutoff. This involves constructing a textual sequence with a missing token representing time-specific information beyond the model’s cutoff. We use each model to tokenize the sentence and predict the missing token. Due to the bi-directional nature of ChronoBERT, the missing token is a masked token in the middle of a sentence, with the surrounding language giving context for the prediction. Consider the following sentence involving an incoming U.S. president, where the president’s name is masked and is the target of the prediction:
Prompt year Accuracy
1993 2001 2009 2017 2021 2025 Pre-cutoff Post-cutoff
Correct output Clinton Bush Obama Trump Biden Trump
BERT Clinton Clinton Obama Obama Obama Obama $2/4$ $0/2$
ModernBERT Clinton Bush Obama Trump Biden Trump $/5$ 5 $/1$ 1
$\text{ChronoBERT}_{1999}$ Roosevelt Hoover Roosevelt Hoover Hoover Washington $0/1$ $0/5$
$\text{ChronoBERT}_{2000}$ Clinton Clinton Clinton Clinton Clinton Wilson $/1$ 1 $0/5$
$\text{ChronoBERT}_{2001}$ Clinton Clinton Clinton Clinton Clinton Washington $1/2$ $0/4$
$\text{ChronoBERT}_{2002}$ Clinton Bush Bush Clinton Bush Clinton $/2$ 2 $0/4$
$\text{ChronoBERT}_{2003}$ Clinton Bush Bush Clinton Bush Clinton $/2$ 2 $0/4$
$\text{ChronoBERT}_{2004}$ Clinton Bush Bush Clinton Bush Clinton $/2$ 2 $0/4$
$\text{ChronoBERT}_{2005}$ Clinton Bush Bush Bush Bush Clinton $/2$ 2 $0/4$
$\text{ChronoBERT}_{2006}$ Clinton Bush Bush Bush Bush Clinton $/2$ 2 $0/4$
$\text{ChronoBERT}_{2007}$ Clinton Bush Bush Bush Bush Monroe $/2$ 2 $0/4$
$\text{ChronoBERT}_{2008}$ Clinton Bush Bush Obama Bush Wilson $/2$ 2 $0/4$
$\text{ChronoBERT}_{2009}$ Clinton Clinton Obama Obama Obama Wilson $2/3$ $0/3$
$\text{ChronoBERT}_{2010}$ Clinton Obama Obama Obama Obama Wilson $2/3$ $0/3$
$\text{ChronoBERT}_{2011}$ Clinton Clinton Obama Obama Obama Wilson $2/3$ $0/3$
$\text{ChronoBERT}_{2012}$ Obama Obama Obama Obama Obama Obama $1/3$ $0/3$
$\text{ChronoBERT}_{2013}$ Clinton Obama Obama Obama Obama Monroe $2/3$ $0/3$
$\text{ChronoBERT}_{2014}$ Clinton Bush Obama Obama Obama Monroe $/3$ 3 $0/3$
$\text{ChronoBERT}_{2015}$ Clinton Clinton Obama Obama Obama Monroe $2/3$ $0/3$
$\text{ChronoBERT}_{2016}$ Clinton Bush Obama Obama Obama Obama $/3$ 3 $0/3$
$\text{ChronoBERT}_{2017}$ Clinton Bush Obama Trump Trump Monroe $/4$ 4 $0/2$
$\text{ChronoBERT}_{2018}$ Clinton Bush Obama Trump Trump Obama $/4$ 4 $0/2$
$\text{ChronoBERT}_{2019}$ Clinton Bush Obama Trump Trump Obama $/4$ 4 $0/2$
$\text{ChronoBERT}_{2020}$ Clinton Bush Obama Trump Trump Trump $/4$ 4 $1/2$
$\text{ChronoBERT}_{2021}$ Clinton Clinton Obama Trump Biden Biden $4/5$ $0/1$
$\text{ChronoBERT}_{2022}$ Clinton Bush Obama Trump Biden Biden $/5$ 5 $0/1$
$\text{ChronoBERT}_{2023}$ Clinton Bush Obama Trump Biden Biden $/5$ 5 $0/1$
$\text{ChronoBERT}_{2024}$ Clinton Bush Obama Trump Biden Biden $/5$ 5 $0/1$
$\text{ChronoBERT}_{1999}$ through $\text{ChronoBERT}_{2024}$ $68/78$ $1/78$
Table 3: Masked-Token Predictions of U.S. Presidents using ChronoBERT
This table displays ChronoBERT’s masked-token predictions in a sentence describing an incoming U.S. president, where the president’s name is the target of the prediction. The input prompt is,